# AROMA: Autonomous Rank-one Matrix Adaptation

### v1

@@ Abstract
As large language models continue to grow in size, parameter-efficient fine-tuning has become increasingly crucial. While low-rank adaptation (LoRA) offers a solution through low-rank updates, its static rank allocation may yield suboptimal results. Adaptive low-rank adaptation (AdaLoRA) improves this with dynamic allocation but remains sensitive to initial and target rank configurations. We introduce AROMA, a framework that automatically constructs layer-specific updates by iteratively building up rank-one components with very few trainable parameters that gradually diminish to zero. Unlike existing methods that employ rank reduction mechanisms, AROMA introduces a dual-loop architecture for rank growth. The inner loop extracts information from each rank-one subspace, while the outer loop determines the number of rank-one subspaces, i.e., the optimal rank. We reset optimizer states to maintain subspace independence. AROMA significantly reduces parameters compared to LoRA and AdaLoRA while achieving superior performance on natural language understanding and commonsense reasoning tasks, offering new insights into adaptive parameter-efficient fine-tuning. The code is available at AROMA.
@@
Hao Nan Sheng1, Zhi-yong Wang1, Mingrui Yang2,3, Hing Cheung So1, 1City University of Hong Kong, 2University of Hong Kong, 3AI Chip Center for Embedded Smart System
@@ 3.1 Adaptive Rank Growth
where $\bm{B}\in\mathbb{R}^{m\times p}$ and $\bm{A}\in\mathbb{R}^{p\times n}$. (4) incorporates a series of rank-one LoRAs. At the beginning of the $p$th outer iteration, a new rank-one LoRA $\bm{b}_{p}\bm{a}_{p}$ is activated for training, while previously calculated $\bm{b}_{1}\bm{a}_{1},\bm{b}_{2}\bm{a}_{2},\cdots,\bm{b}_{p-1}\bm{a}_{p-1}$ are frozen and merged as a single matrix $\bm{B}_{p-1}\bm{A}_{p-1}$.
@@
where $\varepsilon_{\mathrm{in}}$ denotes the inner convergence tolerance, and $\varDelta T_{\mathrm{in}}$ is the inner checking interval. We evaluate (5) every $\varDelta T_{\mathrm{in}}$ steps, and if it is satisfied, the inner loop terminates, and the training of $\bm{b}_{p}\bm{a}_{p}$ is completed.
Once the inner loop ends, we check for outer loop convergence before proceeding to the next outer loop step. Here we use a relative weight change criterion between the $(p-1)$th and the $p$th outer steps defined as:
@@
where $\varepsilon_{\mathrm{out}}$ denotes the outer convergence tolerance. If (6) is satisfied, the outer loop will terminate.
@@ 4.1 Baselines
LoRA-based Methods. 1) LoRA (Hu et al., 2022); 2) AdaLoRA (Zhang et al., 2023a); and 3) ReLoRA (Lialin et al., 2024), which trains $K$ rank-$r$ matrices sequentially and merges them. While ReLoRA is designed for pretraining, it can be regarded as a reduced version of our method, where $T_{\mathrm{in}}$ and $T$ are fixed for all modules, and (5) and (6) are omitted. Therefore, we incorporate it to highlight the effectiveness of AriLoRA’s adaptability and flexibility.
@@ 4.2 Natural Language Understanding
AROMA 1.77M 89.31 83.70 79.12 78.50 83.85 87.43 72.06 94.09 83.51
Table 3: Comparative performance of different fine-tuning schemes for LLaMA3-8B on Commonsense170K dataset. We report accuracy for all tasks. Note that "#Param" reflects the number of initial trainable parameters, and AROMA’s average #Param is even less.
@@ 4.3 Commonsense Reasoning
Training Details. We assign the learning rates $lr$=1E-4, inner tolerance $\varepsilon_{\mathrm{in}}$=0.1, and outer tolerance $\varepsilon_{\mathrm{out}}$=1E-3. We apply AROMA to three weight matrices in the self-attention layer: $\bm{W}_{q}$, $\bm{W}_{k}$, $\bm{W}_{v}$, and two in the FFN: $\bm{W}_{up}$, and $\bm{W}_{down}$. After fine-tuning, the resultant model is evaluated on each of the eight benchmarks in terms of accuracy. Detailed hyperparameter settings are found in Appendix G.2.
@@ G.2 Commonsense Reasoning Task
$\alpha$ 2

### v2

@@ Abstract
As large language models continue to grow in size, parameter-efficient fine-tuning (PEFT) has become increasingly crucial. While low-rank adaptation (LoRA) offers a solution through low-rank updates, its static rank allocation may yield suboptimal results. Adaptive low-rank adaptation (AdaLoRA) improves this with dynamic allocation but remains sensitive to initial and target rank configurations. We introduce AROMA, a framework that automatically constructs layer-specific updates by iteratively building up rank-one components with very few trainable parameters that gradually diminish to zero. Unlike existing methods that employ rank reduction mechanisms, AROMA introduces a dual-loop architecture for rank growth. The inner loop extracts information from each rank-one subspace, while the outer loop determines the number of rank-one subspaces, i.e., the optimal rank. We reset optimizer states to maintain subspace independence. AROMA significantly reduces parameters compared to LoRA and AdaLoRA while achieving superior performance on natural language understanding and commonsense reasoning tasks, offering new insights into adaptive PEFT. The code is available at AROMA.
@@
Hao Nan Sheng1, Zhi-yong Wang1, Mingrui Yang2,3, Hing Cheung So1, 1City University of Hong Kong 2The University of Hong Kong 3AI Chip Center for Embedded Smart System hnsheng2-c@my.cityu.edu.hk
@@ 3.1 Adaptive Rank Growth
where $\bm{B}\in\mathbb{R}^{m\times p}$ and $\bm{A}\in\mathbb{R}^{p\times n}$.
AROMA learns a series of rank-one LoRAs. At the beginning of the $p$th outer iteration, a new rank-one LoRA $\bm{b}_{p}\bm{a}_{p}$ is activated for training, while previously calculated $\bm{b}_{1}\bm{a}_{1},\bm{b}_{2}\bm{a}_{2},\cdots,\bm{b}_{p-1}\bm{a}_{p-1}$ are frozen and merged as a single matrix $\bm{B}_{p-1}\bm{A}_{p-1}$.
@@
where $\varepsilon_{\mathrm{in}}$ denotes the inner convergence tolerance, and $\varDelta T_{\mathrm{in}}$ is the inner checking interval. We evaluate (5) every $\varDelta T_{\mathrm{in}}$ steps, and if it is satisfied, the inner loop terminates, and the training of $\bm{b}_{p}\bm{a}_{p}$, viz., current rank-one LoRA, is completed.
When to stop? Once the inner loop ends, we check for outer loop convergence before proceeding to the next outer loop step. Here we use a relative weight change criterion between the $(p-1)$th and the $p$th outer steps defined as:
@@
where $\varepsilon_{\mathrm{out}}$ denotes the outer convergence tolerance. If (6) is satisfied, the outer loop will terminate, viz., training of $\varDelta\bm{W}$ is completed.
@@ 4.1 Baselines
LoRA-based Methods. 1) LoRA; 2) AdaLoRA; and 3) ReLoRA (Lialin et al., 2024), which trains $K$ rank-$r$ matrices sequentially and merges them. While ReLoRA is designed for pretraining, it can be regarded as a reduced version of our method, where $T_{\mathrm{in}}$ and $T$ are fixed for all modules, and (5) and (6) are omitted. Therefore, we incorporate it to highlight the effectiveness of AriLoRA’s adaptability and flexibility.
@@ 4.2 Natural Language Understanding
AROMA $r$=1 1.77M 89.31 83.70 79.12 78.50 81.85 87.43 71.16 93.79 83.11
AROMA $r$=8 14.16M 89.48 84.79 79.62 78.76 83.98 87.22 73.74 94.36 83.85
Table 3: Comparative performance of different fine-tuning schemes for LLaMA3-8B on Commonsense170K dataset. We report accuracy for all tasks. Results with $\lozenge$" are retrieved from (Liu et al., 2024b). Note that "#Param" reflects the number of initial trainable parameters, and AROMA’s average #Param is even less.
@@ 4.3 Commonsense Reasoning
Training Details. Apart from AROMA under the previous setting, we additionally increase the rank of each LoRA update to 8 to accommodate this complex task. We apply AROMA to three weight matrices in the self-attention layer: $\bm{W}_{q}$, $\bm{W}_{k}$, $\bm{W}_{v}$, and two in the FFN: $\bm{W}_{up}$, and $\bm{W}_{down}$. After fine-tuning, the resultant model is evaluated on each of the eight benchmarks in terms of accuracy. Detailed hyperparameter settings are found in Appendix G.2.
@@ G.2 Commonsense Reasoning Task
Eval Batch Size 8
[…]
$\alpha$ 16
Max Seq. Len. 256
Batch Size 32
Epoch 15
Learning Rate 1E-4
$T$ 80,000
$T_{\mathrm{in}}$ 2000
$\varDelta T_{\mathrm{in}}$ 10
$\varepsilon_{\mathrm{in}}$ 0.1
$\varepsilon_{\mathrm{out}}$ 1E-2